OpenStudy (anonymous):

What is the slope of the line whose equation is 12 = 4x – 6y? ?/?

OpenStudy (haseeb96):

Y=MX+b o x So here m is the slope of the line so just compare your equation with this formula and telel me what is the answer.

OpenStudy (anonymous):

12=4x-6y 6y+12=4x 6y=4x-12 y=(2/3)x-2 m = 2/3 So the slope is 2/3